# Multi-Step Reasoning with Large Language Models, a Survey

### v2

@@ Abstract.
Language models with billions of parameters exhibit in-context learning abilities, enabling few-shot learning on tasks that the model was not specifically trained for. Traditional models achieve breakthrough performance on language tasks, but do not perform well on basic reasoning benchmarks. However, a new in-context learning approach, Chain-of-thought, has demonstrated strong multi-step reasoning abilities on these benchmarks.
The research on LLM reasoning abilities started with the question whether LLMs can solve grade school math word problems, and has expanded to other tasks in the past few years. This paper reviews the field of multi-step reasoning with LLMs. We propose a taxonomy that identifies different ways to generate, evaluate, and control multi-step reasoning. We provide an in-depth coverage of core approaches and open problems, and we propose a research agenda for the near future.
We find that multi-step reasoning approaches have progressed beyond math word problems, and can now successfully solve challenges in logic, combinatorial games, and robotics, sometimes by first generating code that is then executed by external tools. Many studies in multi-step methods are using reinforcement learning for finetuning, external optimization loops, in context reinforcement learning, and self-reflection.
@@ 1. Introduction
Language generation tasks are solved well by LLMs with prompt-based learning. On the other hand tasks that require reasoning, such as grade school math word problems, are more difficult for LLMs (Cobbe et al., 2021). Spurred-on by the impressive performance on language tasks, much research has focused on understanding the reason for the poor performance of LLMs on reasoning tasks, and how it can be improved. Among this research, the Chain-of-thought paper by Wei et al. (2022b) stands out. This work, and later work by Kojima et al. (2022), showed that adding a simple instruction to the prompt, Let’s think step by step, can provoke an LLM to perform the required intermediate reasoning steps to answer difficult problems in a multi-step approach. Subsequently, performance on math word benchmarks has increased markedly. Much of the performance increase of models such as OpenAI o1, o3 (Huang et al., 2024) and DeepSeek R1 (Shao et al., 2024; Guo et al., 2025) is attributed to multi-step reasoning methods as reviewed here, with reinforcement learning playing an increasingly important role, both for in-context learning and finetuning (Xu et al., 2025; Wu et al., 2024a; Du et al., 2025; Chu et al., 2025).
@@ 2. Background, Scope and Selection of Papers
Reasoning has a long and active history in AI, in logical inference, and in other fields, such as commonsense and analogical reasoning (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990). Indeed, some of the early criticism on LLMs was that they could not reason, and that they made obvious and basic reasoning errors, showing a lack of understanding. Berglund et al. (2024) provide the example of a model that is trained to report that “Valentina Tereshkova was the first woman to travel to space,” but is not able to answer the question, “Who was the first woman to travel to space?” pointing indeed to a lack of semantic understanding. Other work suggests that results are less generalizable and transferable than often assumed, showing how base-10 arithmetic skills do not transfer to base-9 arithmetic problems (Wu et al., 2024b). Different kinds of reasoning have played an important role in LLM research.
@@ 2.1. Scope
The interest in Chain-of-though has stimulated more research into language and logic. Logic-LM combines LLMs with symbolic logic solvers in an agentic way, where the symbolic solver is called by the LLM as a tool (Pan et al., 2023). Xu et al. (2024a) use a similar approach. Relevant logic-reasoning benchmarks are PrOntoQA (Saparov and He, 2023), ProofWriter (Tafjord et al., 2020), FOLIO (Han et al., 2022), LogicalDeduction (Srivastava et al., 2022), and AR-LSAT (Zhong et al., 2022). The logic-LLM approaches often combine Chain-of-thought and symbolic solvers in an agentic-LLM approach (Plaat et al., 2025), or transform the problem in an intermediate formal language. The use of LLMs for generating formal languages is further discussed in section 4.4.
More recently, LLMs are also shown to struggle with analogies and analogical reasoning (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990), a topic that is related to self-reflection (see Sections 3.3.3 and 4.3.3).
@@
The question whether LLMs can reason prompted the development of the GSM8k benchmark (Cobbe et al., 2021) in 2021, a benchmark of some of easy reasoning problems (of the type: “Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? Answer: 72”). This benchmark was specifically designed to test if LLMs can answer basic mathematical multi-step reasoning questions, or if they can not reason at all. Initially, performance was poor, but a year later Wei et al. (2022b) showed that by providing multi-step examples in the prompt—a chain of thought—much better performance was possible. Many works followed, and reasoning in language models has become an active area of study. In this survey, we focus on prompt-based multi-step reasoning algorithms. After starting with math word problems, work has more recently been extended to combinatorial games, puzzles, robotics, logic, and other fields of reasoning. At the end of this survey, we discuss connections to other fields, such as self-reflection and in-context reinforcement learning.
@@
Before we discuss the works on reasoning, this section reviews background terminology on LLMs. Our overview is brief. Excellent recent general surveys on LLMs are, for example, Minaee et al. (2024) and Zhao et al. (2023). Other works focus on the nature of reasoning and its definition (Yu et al., 2023; Chu et al., 2024a; Huang and Chang, 2023), on evaluating logical reasoning in LLMs (Mondorf and Plank, 2024; Liu et al., 2023a; Pan et al., 2023), on reinforcement learning in reasoning (Xu et al., 2025), on reasoning in language (Yu et al., 2023), or on Chain-of-thought itself (Chu et al., 2024b).
@@
The papers in this survey were selected as follows. We started by selecting papers on the ability to solve math word benchmarks (as a proxy for reasoning ability), that contained the search terms reasoning and large language model in their title or abstract, with a focus on papers that reference the Chain-of-thought paper. Although multi-step reasoning with LLMs initially was aimed at solving math world problems, it is now wider, including benchmarks and approaches for computer code, game play, puzzles, robot movement, and webpage navigation (see Table 2). We selected recent papers (two years prior to the writing of the survey) that show experimental results on selected benchmark datasets.
@@
We focus on prompt-based, in-context learning, methods based on Chain-of-thought, that are used in reasoning LLMs such as OpenAI o1 and o3 (Xu et al., 2025; Wu et al., 2024a; Huang et al., 2024). We also include papers that work with external algorithms, finetuning or supervised learning that have contributed to the approaches.
@@ 2.2. Language Model Training Pipeline
In training an LLM, the first stage is to acquire a large, general, unlabeled, high-quality text corpus. Some considerations on the selection of the texts are discussed in Brown et al. (2020). The next stage is pretraining the transformer model (Vaswani et al., 2017) on this large corpus. This stage yields a generative language model. Pretraining is done using self-supervised autoregressive training on the unlabeled dataset (text corpus). Then the general model is finetuned to a specific (narrow) task. This can be done using supervised learning with a new labeled dataset consisting of prompts and answers (supervised finetuning, SFT, for example with low-rank optimization (Hu et al., 2022; Wei et al., 2022a; Minaee et al., 2024), or reinforcement learning (Sutton and Barto, 2018; Chu et al., 2025), specific for the task at hand. A small number of the surveyed papers also address the finetuning stage. A specific form of finetuning is instruction tuning, to improve instruction following for a certain task. Instruction tuning is supervised by a labeled dataset of instruction prompts and corresponding outputs. Depending on the purpose of the model the next step is alignment of the finetuned model with user expectations (preference alignment). Alignment is usually performed to ensure that the model produces more ethically and socially acceptable answers, preventing, for example, hate speech. The machine learning method that is commonly used in this stage Reinforcement Learning with Human Feedback (Ouyang et al., 2022) or Direct Preference Optimization (Rafailov et al., 2024). Optionally, model training can be optimized to improve cost-effectiveness, for example, mixed precision training (Micikevicius et al., 2018), quantization (Jacob et al., 2018), or knowledge distillation (Xu et al., 2024b; Gu et al., 2023).
Once the model has been trained in the steps described above, it can be used further in the inference stage. Here, the model is used to provide an answer to the prompt. The inference-time stage is post-training, no model parameters are changed anymore (Dong et al., 2023; Brown et al., 2020); in-context learning, or prompt-learning, takes place in this stage. Due to its low barrier of entry, without the need for expensive training or finetuning, this form of advanced prompt engineering has become a popular training method. This is the stage on which most of the surveyed papers focus, using prompts for the LLM to perform a complex multi-step reasoning task. The following section provides a brief introduction to in-context learning.
@@ 2.3. In-Context Learning
In large models, beyond hundreds of billions of parameters, a new kind of learning has emerged, that has been called in-context learning or prompt-learning (Brown et al., 2020). It occurs not when the model is trained, but when it is used, at inference time. Since no parameters are changed in this stage, it is not a model training stage; in-context-learning “learns” inside the context, or prompt, using information that is already encoded in the trained model parameters and the prompt, not by training the model anymore. In-context learning is often able to give good results with few examples, so-called few-shot learning, learning from the few examples in combination with the knowledge of the model. The large size of the model, containing rich and general knowledge, is enabling the few-shot learning (see Dong et al. (2023) for a survey).
@@
In in-context learning, a prompt, consisting of a piece of demonstration context, is concatenated with a query question, and is given to the language model, for text generation (Liu et al., 2023b). For example, when the task is emotion recognition in a social media post, “I missed the bus today,” can be followed by “I felt so [___]”, and the model could answer with “bad”. Alternatively, for translation, we could follow “I missed the bus today,” by “French: [___]” to request a translation (Liu et al., 2023b). The prompt contains background information that is recognized by the model, selecting the desired model context. In-context learning works when language models contain enough knowledge, allowing them to generalize on the (few) examples provided in the prompt.
@@
Contexts that contain a few examples are said to perform few-shot learning. Contexts that contain only instructions with zero examples are said to perform zero-shot learning. In-context learning takes place at inference time, after the computationally intensive training stages where parameters have been pretrained and finetuned, when the model is queried by the user to provide answers. No parameters are changed anymore with in-context learning. This is quite different from the common approach in supervised deep learning—or self-supervised deep learning—where large datasets are used during training to update model parameters with backward propagation in lengthy and costly training epochs (Goodfellow et al., 2016). Indeed, in-context learning takes place fully at inference time, no parameters are trained, instead, learning now refers to adjusting the answers to the examples in the prompt and the internal knowledge acquired during training. Common approaches to few-shot learning, such as metalearning, do include training and finetuning of parameters to achieve generalization, and are computationally expensive (see, for example, (Finn et al., 2017) or (Huisman et al., 2021; Hospedales et al., 2021) for a survey). In-context learning, in comparison, is computationally cheap, and it has become a popular research approach..
@@
Prompts provide a user-friendly interface to LLMs. The success of in-context learning tends to be quite sensitive to the way a prompt is formulated; a new field called prompt engineering has emerged to help human users learn how to make LLMs do what we want them to do (Radford et al., 2019; Wei et al., 2022a; Giray, 2023; Sahoo et al., 2024). The current survey thus discusses advanced prompt engineering methods.
@@ 2.4. Multi-step Reasoning Benchmarks
It can be difficult to directly compare different benchmarking results, since there may be differences in the way the experiments were conducted. Kamoi et al. (2024) provide a critical analysis of benchmarking problems in LLM reasoning. Despite these challenges, general conclusions can be drawn. Section 4.1 provides guidance on how to choose reasoning approaches for different problem types.
@@ 3. Step Generation, Evaluation and Control
This survey examines how LLMs based on the transformer architecture can be prompted to solve multi-step reasoning tasks. The Chain-of-thought paper shows how a simple command could prompt an LLM to perform reasoning steps, yielding much better performance in math word problems. Since then much research has further explored this approach, trying to build stronger general problem solvers for other types of reasoning problems.
@@
    Generate: prompt the model for the generation of steps,
@@
    Evaluate: prompt for evaluation of the generated steps,
@@
    Control: prompt for control of the number of steps that are generated and how deep ahead the reasoning process will look.
This three-stage pipeline is the basis of our taxonomy. We will now discuss the three stages more deeply; for ease of reference they are numbered according to the Subsection in which they are described in more detail (3.1, 3.2, 3.3). Please also refer to Figure 1, or Table 2, for a diagram of the categories and subcategories of different approaches for the generation, evaluation, and control of reasoning steps. [^1] We show the approaches in the Figure in their main category only. Some approaches show innovations in two categories, and are shown twice. (Since all approaches have a generation, an evaluation, and a control aspect, all could in principle occur three times, and all three columns can be found in Table 2).
@@
The first stage is to create a prompt that instructs the LLM to generate reasoning steps. The problem must be split into substeps. This can be achieved with a problem-specific prompt that contains elements of the problem, such as: “First calculate how many marbles Mary had originally, then how many her friend had, and finally how many they had together.” In general, it is possible to prompt an LLM to fill in the blanks in a step-by-step fashion. In the papers that we discuss, there are three main approaches for generating the step-by-step prompt, numbered with the Subsection in which the approaches are described. First, the prompt may be handcrafted for the problem by the researchers: (3.1.1) hand-written prompt. Second, the prompt or prompts may come from a source that is external to the model, such as another model or dataset: (3.1.2) external knowledge-based prompt. Third, the model itself can be prompted to generate a (series of) prompt(s) to analyze the problem (3.1.3) model-generated prompt. As we will see, all three approaches have their advantages and disadvantages.
@@
After the prompt has been generated and the model has answered it, the next stage is to evaluate the quality of this answer. Such evaluation is often necessary to improve and perform well on the benchmark. Again, we see three main approaches for substep evaluation. First, the steps may be evaluated by the model itself: (3.2.1) self-assessment. Second, an external program can be used to evaluate the steps. For example, when the steps are expressed as computer code, an external interpreter or compiler can be used to check the validity and the outcome: (3.2.2) tool-based evaluation. Finally, an external model can be used, LLM or otherwise. For example, in robotics, an external physics model can determine if certain actions are physically possible: (3.2.3) external model validation.
@@
Again, we distinguish three approaches. Most papers use a (3.3.1) greedy selection approach: a single prompt with a single chain of steps is generated, and these steps are directly executed and followed. The second approach is to generate an (3.3.2) ensemble strategy of reasoning steps, evaluate them, combine the individual results, and present them as the result of the ensemble. Finally, a full tree-search or a (3.3.3) reinforcement learning (RL) algorithm can be used as scaffolding. In this case, when a step is followed and evaluated, the LLM can roll back and try a different reasoning step (Plaat, 2020). Going further, a full reinforcement learning approach can be used (Sutton and Barto, 2018; Plaat, 2022) to find an optimal policy for the sequential decision process. In general a Markov Decision Process of state, action, transition, and reward function can be specified, and step control can become a process where prompts are generated dynamically under the control of an external RL algorithm, or as in-context reinforcement learning (ICRL) (Lee et al., 2023).

### v3

@@ Abstract.
Large language models (LLMs) with billions of parameters exhibit in-context learning abilities, enabling few-shot learning on tasks that the model was not specifically trained for. Traditional models achieve breakthrough performance on language tasks, but do not perform well on basic reasoning benchmarks. However, a new in-context learning approach, Chain-of-thought, has demonstrated strong multi-step reasoning abilities on these benchmarks.
The research on LLM reasoning abilities started with the question whether LLMs can solve grade school math word problems, and has expanded to other tasks in the past few years. This article reviews the field of multi-step reasoning with LLMs. We propose a taxonomy that identifies different ways to generate, evaluate, and control multi-step reasoning. We provide an in-depth coverage of core approaches and open problems, and we propose a research agenda for the near future.
We find that multi-step reasoning approaches have progressed beyond math word problems, and can now successfully solve challenges in logic, combinatorial games, and robotics, sometimes by first generating code that is then executed by external tools. Many studies in multi-step methods use reinforcement learning for finetuning, external optimization loops, in-context reinforcement learning, and self-reflection.
@@ 1. Introduction
Language generation tasks are solved well by LLMs with prompt-based learning. On the other hand, tasks that require reasoning, such as grade school math word problems, are more difficult for LLMs (Cobbe et al., 2021). Spurred-on by the impressive performance on language tasks, much research has focused on understanding the reason for the poor performance of LLMs on reasoning tasks, and how it can be improved. Among this research, the Chain-of-thought paper by Wei et al. (2022b) stands out. This work, and later work by Kojima et al. (2022), showed that adding a simple instruction to the prompt, Let’s think step by step, can provoke an LLM to perform the required intermediate reasoning steps to answer difficult problems in a multi-step approach. Subsequently, performance on math word benchmarks has increased markedly. Much of the performance increase of models such as OpenAI o1, o3 (Huang et al., 2024) and DeepSeek R1 (Shao et al., 2024; Guo et al., 2025) is attributed to multi-step reasoning methods as reviewed here, with reinforcement learning playing an increasingly important role, both for in-context learning and finetuning (Xu et al., 2025; Wu et al., 2024a; Du et al., 2025; Chu et al., 2025).
@@ 2. Background, Scope and Selection of Papers
Reasoning has a long and active history in AI, in logical inference, and in other fields, such as commonsense and analogical reasoning (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990). Indeed, some of the early criticism on LLMs was that they could not reason, that they made obvious and basic reasoning errors, showing a lack of understanding. Berglund et al. (2024) provide the example of a model that is trained to report that “Valentina Tereshkova was the first woman to travel to space,” but is not able to answer the question, “Who was the first woman to travel to space?” pointing to a lack of semantic understanding. Other work suggests that results are less generalizable and transferable than often assumed, showing how base-10 arithmetic skills do not transfer to base-9 arithmetic problems (Wu et al., 2024b). More recently, LLMs are also shown to struggle with analogies and analogical reasoning (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990) (see Sections 3.3.3 and 4.3.3). Different kinds of reasoning play an important role in LLM research.
@@ 2.1. Scope
The question whether LLMs can reason prompted the development of the GSM8k benchmark (Cobbe et al., 2021) in 2021. GSM8k is a benchmark of 8500 easy reasoning problems (of the type: “Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May. How many clips did Natalia sell altogether in April and May? Answer: 72”). This benchmark was specifically designed to test if LLMs can answer basic mathematical multi-step reasoning questions, or if they can not reason at all. Initially, performance was poor, but a year later Wei et al. (2022b) showed that by providing multi-step examples in the prompt—a chain of thought—much better performance was possible. Many works followed, and reasoning in language models has become an active area of study. In this survey, we focus on prompt-based multi-step reasoning algorithms. After starting with math word problems, work has more recently been extended to combinatorial games, puzzles, robotics, logic, and other fields of reasoning. At the end of this survey, we discuss connections to other fields, such as self-reflection and in-context reinforcement learning.
@@
Before we discuss the works on reasoning, this section reviews background terminology on LLMs. Our overview is brief and we focus on the multi-step reasoning approaches that were inspired by Chain-of-thought. Excellent related surveys exist. For a longer overview of general LLM topics, see, for example, the surveys Minaee et al. (2024) and Zhao et al. (2023). For surveys that focus on the nature of reasoning and its definition, see (Yu et al., 2023; Chu et al., 2024a; Huang and Chang, 2023). For surveys on evaluating logical reasoning in LLMs, see (Mondorf and Plank, 2024; Liu et al., 2023a; Pan et al., 2023). For a survey on reinforcement learning in reasoning, see (Xu et al., 2025), for reasoning in language see (Yu et al., 2023), and for a survey on Chain-of-thought itself, see (Chu et al., 2024b).
@@
The papers for this survey were selected as follows. We started by selecting papers on the ability to solve math word benchmarks (as a proxy for reasoning ability), that contained the search terms reasoning and large language model in their title or abstract, with a focus on papers that reference the Chain-of-thought paper. Although multi-step reasoning with LLMs initially was aimed at solving math world problems, it is now wider, including benchmarks and approaches for computer code, game play, puzzles, robot movement, and webpage navigation (see Table 2). We selected recent papers (two years prior to the writing of the survey) that show experimental results on selected benchmark datasets.
@@
We focus on prompt-based, in-context learning, methods based on Chain-of-thought, that are used in reasoning LLMs such as OpenAI o1 and o3 (Xu et al., 2025; Wu et al., 2024a; Huang et al., 2024). We include a few papers that work with external algorithms, finetuning, or supervised learning, that have contributed significantly to the in-context approaches.
@@ 2.2. Language Model Training Pipeline
In training an LLM, the first stage is to acquire a large, general, unlabeled, high-quality text corpus. Some considerations on the selection of the texts are discussed in Brown et al. (2020). The next stage is pretraining the transformer model (Vaswani et al., 2017) on this large corpus. This stage yields a generative language model. Pretraining is done using self-supervised autoregressive training on the unlabeled dataset (text corpus). Then the general model is finetuned to a specific (narrow) task, for example, question answering. This can be done using supervised learning with a new labeled dataset consisting of prompts and answers (supervised finetuning, SFT), for example with low-rank optimization (Hu et al., 2022; Wei et al., 2022a; Minaee et al., 2024), or reinforcement learning (Sutton and Barto, 2018; Chu et al., 2025). A specific form of finetuning is instruction tuning, to improve instruction following for a certain task. Instruction tuning is supervised by a labeled dataset of instruction prompts and corresponding outputs. Depending on the purpose of the model the next step is alignment of the finetuned model with user expectations (preference alignment). Alignment is usually performed to ensure that the model produces more ethically and socially acceptable answers, preventing, for example, hate speech. A machine learning method that is commonly used in this stage is Reinforcement Learning with Human Feedback (Ouyang et al., 2022), Direct Preference Optimization (Rafailov et al., 2024), or Reinforcement Learning with Verifiable Rewards (Lambert et al., 2024). Optionally, model training can be optimized to improve cost-effectiveness, for example, mixed precision training (Micikevicius et al., 2018), quantization (Jacob et al., 2018), or knowledge distillation (Xu et al., 2024b; Gu et al., 2023).
Once the model has been trained in the steps described above, it can be used further in the inference stage. Here, the model is used to provide an answer to the prompt. The inference-time stage is post-training, no model parameters are changed anymore (Dong et al., 2023; Brown et al., 2020); in-context learning, or prompt-learning, takes place in this stage.
In-context learning has a low barrier of entry. Since there is no need for expensive pretraining or finetuning, this form of advanced prompt engineering has become a popular training method. This is the stage on which most of the surveyed papers focus, using prompts for the LLM to perform a complex multi-step reasoning task. The following section provides a brief introduction to in-context learning.
@@ 2.3. In-Context Learning
In large models, beyond hundreds of billions of parameters, a new kind of machine learning has emerged, that has been called in-context learning or prompt-learning (Brown et al., 2020). It occurs not when the model’s parameters are trained, but when the model is used, at inference time. Since no parameters are changed at this stage, it is not a model training stage; in-context-learning “learns” inside the context, or prompt, using information that is already encoded in the trained model parameters and the prompt, not by training the model anymore. In-context learning is often able to give good results with few examples, so-called few-shot learning, learning from the few examples in the prompt in combination with the knowledge of the model. The large size of the model, containing rich and general knowledge, is enabling the few-shot learning (see Dong et al. (2023) for a survey).
@@
In in-context learning, a prompt, consisting of a piece of demonstration context, is concatenated with a query question, and is given to the language decoder model, for text generation (Liu et al., 2023b). For example, when the task is emotion recognition in a social media post, “I missed the bus today,” can be followed by “I felt so [___]”, and the model could answer with “bad”. Alternatively, for translation, we could follow “I missed the bus today,” by “French: [___]” to request a translation (Liu et al., 2023b). The prompt contains background information that is recognized by the model, selecting the desired model context. In-context learning works when language models contain enough knowledge, allowing them to generalize on the (few) examples provided in the prompt.
@@
Contexts that contain a few examples are said to perform few-shot learning. Contexts that contain only an instruction, with zero examples, are said to perform zero-shot learning. In-context learning takes place at inference time, after the computationally intensive training stages where parameters have been pretrained and finetuned, when the model is queried by the user to provide answers. No parameters are changed anymore with in-context learning. This is quite different from the common approach in supervised deep learning—or self-supervised deep learning—where large datasets are used during training to update model parameters with backward propagation in lengthy and costly training epochs (Goodfellow et al., 2016). Indeed, in-context learning takes place fully at inference time, no parameters are trained, instead, learning now refers to adjusting the answers to the examples in the prompt and the internal knowledge acquired during training. Common approaches to few-shot learning, such as metalearning, do include training and finetuning of parameters to achieve generalization, and are computationally expensive (see, for example, (Finn et al., 2017) or (Huisman et al., 2021; Hospedales et al., 2021) for a survey). In-context learning, in comparison, is computationally cheap, and has become a popular research approach.
@@
Prompts provide a user-friendly interface to LLMs. The success of in-context learning tends to be quite sensitive to the way in which a prompt is formulated; a new field called prompt engineering has emerged to help human users to learn how to make LLMs do what they want them to do (Radford et al., 2019; Wei et al., 2022a; Giray, 2023; Sahoo et al., 2024). The current survey thus discusses advanced prompt engineering methods.
@@ 2.4. Multi-step Reasoning Benchmarks
In this survey we will mention benchmark results as reported by the surveyed papers, for the sake of concreteness. However, it can be difficult to directly compare different benchmarking results, since there may be differences in the way the experiments were conducted. Kamoi et al. (2024) provide a critical analysis of benchmarking problems in LLM reasoning. Despite these challenges, general conclusions can be drawn. Section 4.1 provides guidance on how to choose reasoning approaches for different problem types.
@@ 3. Step Generation, Evaluation and Control
This survey examines how LLMs based on the transformer architecture can be prompted to solve multi-step reasoning tasks. The Chain-of-thought paper shows how a simple command can prompt an LLM to perform reasoning steps, yielding much better performance in math word problems. Much research has further explored this approach, trying to build stronger general problem solvers for other types of reasoning problems.
@@
    Generate: prompt the model to generate reasoning steps,
@@
    Evaluate: prompt the model to evaluate the generated steps,
@@
    Control: prompt the model to control the number of steps that are generated and how deep ahead the reasoning process will look.
This three-stage pipeline is the basis of our taxonomy. We will now discuss the three stages more deeply; for ease of reference they are numbered according to the Subsection in which they are described in more detail (3.1, 3.2, 3.3). Please also refer to Figure 1, and to Table 2, for a diagram of the categories and subcategories of different approaches for the generation, evaluation, and control of reasoning steps. [^1] We show the approaches in the Figure in their main category only. Some approaches show innovations in two categories, and are shown twice. (Since all approaches have a generation, an evaluation, and a control aspect, all could in principle occur three times, and all three columns can be found in Table 2).
1. Generation2. Evaluation3. Control Hand-written prompt Model-generated prompt Prompt using external knowledge Scratchpad; Nye et al. (2022) [supervised]
Chain-of-Thought; Wei et al. (2022b)
ZS-CoT; Kojima et al. (2022)
Auto-CoT; Zhang et al. (2023)]
Complexity; Fu et al. (2022)
Buffer-of-Thoughts; Yang et al. (2024)
Self-Ask; Press et al. (2023) Self-Assessment Tool-based Evaluation External Model Validation Self-Verification; Weng et al. (2023)
Self-Consistency; Wang et al. (2023c)
Codex; Chen et al. (2021)
Self-Debugging; Chen et al. (2024)
Fun-Search; Romera-Paredes et al. (2024)
LLaMea; van Stein and Bäck (2024)
MathPrompter; Imani et al. (2023)
Program-of-Thoughts; Chen et al. (2023)
Program-Aided-Language; Gao et al. (2023)
Finetuning:
  Refiner; Paul et al. (2024a)
  Self-Corrector; Welleck et al. (2023)
  Self-Improvement; Huang et al. (2023)
Robot:
  SayCan; Ahn et al. (2022)
  Inner Monologue; Huang et al. (2022)
Data Augmentation:
Self-Taught Reasoner; Zelikman et al. (2022)
Greedy Selection Ensemble Strategy Reinforcement Learning Complexity (Fu et al. 2022)
Least-to-Most (Zhou et al. 2022)
Self-Consistency; Wang et al. (2023c)
Self-Verification; Weng et al. (2023)
MathPrompter; Imani et al. (2023)
Program-Aided-Language; Gao et al. (2023)
Progressive-Hint; Zheng et al. (2023)
Self-Refine; Madaan et al. (2023)
Tree-of-Thoughts; Yao et al. (2024)
Buffer-of-Thoughts; Yang et al. (2024)
Algorithm-of-thoughts; Sel et al. (2023)
Beam-Search; Xie et al. (2024)
ReAct; Yao et al. (2023)
Reflexion; Shinn et al. (2024)
Voyager; Wang et al. (2024a)
@@
The first stage is to create a prompt that instructs the LLM to generate reasoning steps. The problem must be split into substeps. This can be achieved with a problem-specific prompt that contains elements of the problem, such as: “First calculate how many marbles Mary had originally, then how many her friend has, and finally how many they have together.” In general, it is possible to prompt an LLM to fill in the blanks in a step-by-step fashion. In the papers that we discuss, there are three main approaches for generating the step-by-step prompt, numbered with the Subsection in which the approaches are described. First, the prompt may be handcrafted for the problem by the researchers: (3.1.1) hand-written prompt. Second, the prompt or prompts may come from a source that is external to the model, such as another model or dataset: (3.1.2) external knowledge-based prompt. Third, the model itself can be prompted to generate a (series of) prompt(s) to analyze the problem (3.1.3) model-generated prompt. As we will see, all three approaches have their advantages and disadvantages.
@@
After the prompt has been generated and the model has answered it, the next stage is to evaluate the quality of this answer. Such evaluation is often necessary to improve the answer and perform well on the benchmark. Again, we see three main approaches for substep evaluation. First, the steps may be evaluated by the model itself: (3.2.1) self-assessment. Second, an external program can be used to evaluate the steps. For example, when the steps are expressed as computer code, an external interpreter or compiler can be used to check the validity of the outcome: (3.2.2) tool-based evaluation. Finally, an external model can be used, LLM or otherwise. For example, in robotics, an external physics model can determine if certain actions are physically possible: (3.2.3) external model validation.
@@
Again, we distinguish three approaches. Most papers use a (3.3.1) greedy selection approach: a single prompt with a single chain of steps is generated, and these steps are directly executed and followed. The second approach is to generate an (3.3.2) ensemble strategy of reasoning steps, evaluate them, combine the individual results, and present them as the result of the ensemble. Finally, a full tree-search or a (3.3.3) reinforcement learning (RL) algorithm can be used as scaffolding (Plaat, 2020). In this case, when a step is followed and evaluated, the LLM can roll back and try a different reasoning step. Going further, a full reinforcement learning approach can be used (Sutton and Barto, 2018; Plaat, 2022) to find an optimal policy for the sequential decision process. In general a Markov Decision Process of state, action, transition, and reward function can be specified, and step control can become a process where prompts are generated dynamically under the control of an external RL algorithm, or as in-context reinforcement learning (ICRL) (Lee et al., 2023; Moeini et al., 2025).